Italian Queue 2014 Prev Next

Jun 18-19

Observer: E. Sani, A. Marchetti, M. Fumana
Telescope Operator: D. Gonzales-Huerta
Support Astronomer: D. Thompson on call
Instrument: MODS/LUCI

Summary

At the beginning of the nigh we have clear sky all over the orison, now wind and low humidity.
The average seeing at the beginning of scientific operations is 1.0 on the GCS.
Thin cirrus passing during the 1st LUCI integration.
With LUCI we got a flexure compensation error and the WFS was not working for few minutes thus an offset failed during acquisition.
AGw got stuck imaging acquisition.
The readout problem persist and get worse at the end of the night.
We also got condensation on the camera

Data Summary

Project_RUN Objects Filter ExpTime Notes
Bellazzini_46_DDT HVC274.68+74 red gradint 1h within the specs PROGRAM COMPLETED
Standard HIP82884 210_zJHK, J-band   thin cirrus, to be repeated
Magrini_24 7C1756 Mask1 210_zJHK, J-band 150min LUCI bias+pickup noise
Cresci_8 campoB Ks 26m LUCI bias+pickup noise. Bad WFS corrections, too close to the Moon

Details

02:45 the TO opens the DOME

03:32 acqMODS HVC274.68+74_LS_recover.acq 1r 0001

03:39 We slew to Bellazzini_46_DDT
  • execMODS HVC274.68+74_LS_3.obs 1r 0002-0003
  • the initial seeing is 1.0 and 1.2 at the end of integration

04:43 execMODS HVC274.68+74_LS_1 1r 0004
  • at the beginning of integration the average seeing is 1.3.
  • The seeing increases during the integration reaching 1.5

05:16 We ask the TO to switch to LUCI

05.36 We are authorized to use LUCI and switch to Magrini_24
  • run the standard star acquisition executeLUCIscript.sh 7C1756Mask1_MOS_210_HeIJ_tell.acq luci.20140619.006-009
  • thin cirrus passing

05:51 executeLUCIscript.sh 7C1756Mask1_MOS_210_J_tell.spec luci.20140619.0010-0015

05:55 executeLUCIscript.sh 7C1756Mask1_MOS_210_H.acq luci.20140619.0016
  • error sending an offset to the telescope (happened after). It seems an instrument problem. The TO calls the SA. We sent again the acq script
  • the GCS stopped while the offset was sent... the TO asks to exit the script. The guider wavefront sensing was not responding, that's why the system exited the script the first time the problem happened.

06:10 We run again the script (the 1st offset is commented and the acquisition command modified: no sky image repeated) luci.20140619.0016-0019
  • Flexure compensation error after the through slit image. We switch the set fex on and off and it worked.

6:28 executeLUCIscript.sh 7C1756Mask1_MOS_210_J_1.spec luci.20140619.0020-0029
  • at the beginning of the integration the seeing is 1.5 on the GCS
  • thin cirrus passing. at 7th offset the seeing is 1.1
  • thin cirrus where passing with a massimum flux variation of 0.25 mag (from the guider)

07:29 executeLUCIscript.sh 7C1756Mask1_MOS_210_J_2.spec luci.20140619.0030-0039
  • at the beginning of the integration the seeing is 0.9 on the GCS and the sky is now clear.
  • at the end of integration the average seeing is 0.9

08:20 executeLUCIscript.sh 7C1756Mask1_MOS_210_J_3.spec luci.20140619.0040-0049
  • at the beginning seeing 1.0 on the GCS, clear sky.
  • thin cirrus passing at the end of the integration. The maximum seeing was 1.2 (a couple of offsets)

09:17 We slew to the standard star
  • executeLUCIscript.sh 7C1756Mask1_MOS_210_HeIJ_tell.acq luci.20140619.0050-0052

09:26 executeLUCIscript.sh 7C1756Mask1_MOS_210_J_tell.spec luci.20140619.0053-0058
  • good telluric

09:32 We slew to Cresci_8 imaging of CampoB
  • executeLUCIscript.sh CampoB _Ks_1.img
  • We got an error the camera weel, we click INITIALIZE and it worked
  • The AGw is stuck

09:51 try to send the preset again
  • executeLUCIscript.sh CampoB _Ks_1.img luci.20140619.0059-0084
  • the target is extremely close to the Moon, the WFS sees the star like R=17.5 mag rather than 14.2. This implies a not so good collimation (RMS 400-1500)
  • Condensation on the camera seen in some frames. The TO calls John Morris

10:47 We run test images to monitor the condensation features luci.20140619.0085-0090

11.02 We run sky flats for program Beltran_01 in the H2 narrowband filter
  • luci.20140619.0091-0100 without dithering, no good
  • executeLUCIscript.sh dither_Ks.flat luci.20140619.0101-110 good off-flats
  • executeLUCIscript.sh dither_Ks.flat luci.20140619.0111-0124 no good, low counts

11:56 the TO decide to close the DOME, intence light is too risky.
Topic revision: r7 - 23 Jun 2014, EleonoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback